Circular No. 6123

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994ak IN NGC 2782 M. W. Richmond, Princeton University; and R. R. Treffers, S. D. Van

Dyk, and A. V. Filippenko, University of California at Berkeley, confirm that there is a new starlike object in NGC 2782, presumably a supernova, close to the position reported on IAUC 6122. Observations were conducted as part of the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. The supernova (R = 17.4 ± 0.8) is located ∼ 58′′ west and 33′′ south of the galaxy’s nucleus in an image obtained on 1994 Dec. 30 UT; it is also barely visible in a mediocre image obtained on Dec. 23, but not visible (upper limit R = 19) in a good image on Dec. 1.

S. Nakano, Sumoto, Japan, forwards the following additional unfiltered CCD magnitudes of SN 1994ak obtained with a 0.25-m Schmidt-Cassegrain by Y. Kushida at Yatsugatake South Base Observatory: 1994 Dec. 25.780 UT, 15.9; 28.591, 16.0; 29.619, 16.0; 31.616, 16.1; 1995 Jan. 1.645, 16.3; 2.600, 16.3.

D. Balam, Climenhaga Observatory, Victoria, provides the following magnitude and precise 2000.0 position for SN 1994ak obtained on Jan. 3.44 UT: V = 18, α = 9^{h}14^{m}01^{s}.47, δ = +40^{o}06′21′′.5; the supernova is indicated as being ∼ 43′′ west and 27′′ south of the galaxy’s center.

SUPERNOVA 1994ai IN NGC 908 M. Cavagna and E. Galliani, Sormano, Italy, report the following pre-

cise position for SN 1994ai, obtained on Jan. 1.80 UT: α = 2^{h}23^{m}06^{s}.17, δ = −21^{o}13′58′′.3 (equinox 2000.0).

V1251 CYGNI This cataclysmic variable (cf. IAUC 5377) is evidently in outburst, as

indicated by the following visual magnitude estimates: 1994 Dec. 22.736 UT, [13.6 (P. Schmeer, Bischmisheim, Germany); 30.756, 13.3: (Schmeer); 31.713, 13.4 (L. Szentaskó, Veresegyház, Hungary); 31.833, 13.2 (Schmeer); 1995 Jan. 1.771, 13.7 (Schmeer). The following precise position for V1251 Cyg has been measured by B. Manning, Stakenbridge, England: α = 21^{h}40^{m}54^{s}.36, δ = +48^{o}39′43′′.1 (equinox 2000.0).

1995 January 4 Daniel W. E. Green